OpenStudy (anonymous):

Given a rectangular prism with dimensions w = 3, l = 4, and h = 6. If you created a second rectangular prism with the length doubled but the height halved (and the width stays the same), which would be the relation of the second volume to the first volume?

OpenStudy (cherylim23):

V1 = 3*4*6 = 72 V2= 2L*H/2*W = L*H*W = 3 x 4 x 6 = 72 Wouldn't they be equal?